OpenStudy (anonymous):

find the derivative of ln(cos^2(e^x))

zepdrix (zepdrix):

We can make this a smidge easier by applying a rule of logarithms:\[\Large \frac{d}{dx}\ln\left(\cos^2e^x\right)\quad=\quad \frac{d}{dx}2\ln\left(\cos e^x\right)\]

zepdrix (zepdrix):

Do you recall the derivative of \(\Large \ln x\) ?

OpenStudy (anonymous):

\[\frac{ 1 }{ x }\] is the derivative of

OpenStudy (loser66):

@zepdrix continue, please. you have using color skill in explaining this stuff. Use it, please

OpenStudy (loser66):

sorry friend, he is sleeping. My method is not as good as him, please try all your best to understand. If you cannot, you have to wait for him. hehehe

zepdrix (zepdrix):

lol :3

OpenStudy (loser66):

XD